1. A person building a brick making machine wants to measure the coefficient of static friction between brick and wood. To do this, she places a 2.00 kg brick on a flat piece of wood, and gradually applies more and more force until the brick moves. She finds that the brick moves when exactly 11.8 N of force is applied. What is the coefficient of static friction?

2. A woman is skiing straight down the snow covered hill. The coefficient of kinetic friction between the skis and the snow is 0.0800. The hill is at 60.0 degree angle form the horizontal. The skier mass is 55.00 kg. What is the magnitude of the force of kinetic friction, and what is the net force along the skiers motion?
1
Expert's answer
2018-11-12T15:55:09-0500

F=μmg

The coefficient of static friction:

μ=F/mg=11.8/((2)(9.81))= 0.601


2. A woman is skiing straight down the snow covered hill. The coefficient of kinetic friction between the skis and the snow is 0.0800. The hill is at 60.0 degree angle form the horizontal. The skier mass is 55.00 kg. What is the magnitude of the force of kinetic friction, and what is the net force along the skiers motion?

Solution

The magnitude of the force of kinetic friction

F=μmg=(0.08)(55)(9.81)=43.2 N.

The net force along the skiers motion

F_net=mg sin⁡60-μmg cos⁡60

F_net=(55)(9.81)(sin⁡60-0.08 cos⁡60 )=446 N.